ThisGirlPretty:

Solve for x. 4x + 8 = 7.2 + 5x

AshtynThatDrummerGirl:

I got this dB^| Give me a moment XD

ThisGirlPretty:

Sure thing XD, Trying to understand how to do this LOL

AshtynThatDrummerGirl:

I'll write the instructions out for you ^.^

AshtynThatDrummerGirl:

4x + 8 = 7.2 + 5x take away 8 from both sides of the equation to get this: 4x = -0.8 + 5x subtract 5x from both sides -x = -0.8 divide both sides by -1 to get x alone x = 0.8 Let me know if you need me to draw it ^.^